OpenStudy (anonymous):

Find the slope of the line using the two plotted points in the graph. (-6,5)(9,-5) 9-(-6)=3 -5-5=-10 3/-10 dont know if this is right

OpenStudy (katrinakaif):

Slope should be -2/3

OpenStudy (katrinakaif):

First use the formula (to find the slope) : y2-y1/ x2-x1 -5 - 5 = -10 over 9 + 6 (Because - times a negative =s a positive 9-(-6) = 9+ 6 giving -10/15 = -2/3

OpenStudy (anonymous):

can you show me how you got that

OpenStudy (katrinakaif):

Do you understand? If so please 'Good Answer' it =) Thankyou!